OpenStudy (anonymous):

Use an appropriate Half-Angle Formula to find the exact value of the expression. cos 112.5°

OpenStudy (turingtest):

\[\cos (112.5°)=\cos ({225°\over2})\]\[\cos^2({\theta \over2})={1\over2}(1+\cos \theta)\]between these two formulas lies your answer...

OpenStudy (anonymous):

I don't get it!